# Co-Training Vision Language Models for Remote Sensing Multi-task Learning

### v1

@@ IV-D1 Benchmark, Metric, and Comparison Setting
We selected the most widely-used aerial image object detection benchmark, DOTA-v1.0 [dota], for our evaluation. The whole DOTA-v1.0 dataset comprises 2,806 high-resolution aerial images and 188,282 object instances across 15 common categories. The proportions of testing set is 1/3. These images were collected from multiple sensors and platforms, and each instance is annotated with a 8 degrees-of-freedom oriented bounding box, capturing the wide variations in object scale, shape, and orientation typical of aerial imagery.
We adopt the Average Precision with no confidence ($\text{AP}_{\text{nc}}$) and report three specific variants: $\text{AP}_{\text{nc}50}$ (IoU threshold is 0.50), $\text{AP}_{\text{nc}75}$ (IoU threshold is 0.75), and $\text{AP}_{\text{nc}50:95}$ (the average $\text{AP}_{\text{nc}}$ computed over IoU thresholds from 0.50 to 0.95 at increments of 0.05). The evaluation is based on the standard MMRotate [mmrotate] evaluation procedure. And the splitting length is set 512 with an overlap of 100.
The conventional object detection baselines are trained using the latest MMRotate [mmrotate], and the details necessary for reproducibility are also provided in the released code. We obtain a reasonable $\text{AP}_{\text{nc}}$ for comparison methods using the following procedure: We first select a threshold for confidence scores to filter out low-score predictions, and then randomize (or, set to 1) the remaining prediction scores. The AP computed under this condition is denoted as $\text{AP}_{\text{nc}}$ of the conventional detector. To determine an appropriate threshold for each detector, we evaluate $\text{AP}_{\text{nc}}$ on the validation set by varying the confidence threshold from 0.00 to 0.95 in increments of 0.05, and select the threshold that yields the highest $\text{AP}_{\text{nc}}$ for subsequent evaluation on the test set.
@@ IV-D2 Results
To the best of our knowledge, only two existing RS VLMs, LMMRotate [lmmrotate] (our conference version) and Falcon [falcon], are capable of performing aerial object detection effectively. However, their common foundation model, Florence-2 [florence2], employs a fixed and relatively large input size of $\times1024$, which already exceeds the input limit of RSCoVLM. Moreover, LMMRotate is trained specifically for detection, while Falcon performs well only on its training set and does not report test results. In addition, Falcon requires multiple inferences per image, making separate predictions for each category, which results in extremely high computational cost. Therefore, we consider RSCoVLM to be the only vision-language model capable of performing multiple tasks while achieving detection performance that is fairly comparable to specialized object detection models, currently.
@@ IV-E1 Benchmark and Metric
We evaluate our model on five standard remote‑sensing scene‑classification benchmarks. The AID [aid] dataset comprises approximately 10,000 images of size $\times600$ pixels across 30 classes. The UCMerced [ucm] dataset consists of 2,100 images of size $\times256$ pixels covering 21 classes. The NWPU‑RESISC45 [resisc] dataset contains 31,500 images of size $\times256$ pixels across 45 classes, with large variation in resolution and scene complexity. The WHU‑RS19 [whurs19] dataset includes around 1,000 high‑resolution patches of size $\times600$ pixels spanning 19 classes. The METER‑ML [meterml] benchmark offers a large‑scale multi‑sensor setup with varied image sizes for extended generalisation evaluation. Together these benchmarks allow a robust assessment of our model’s generalisation across dataset scale, class‑set size, imaging conditions and spatial resolutions.
We report overall accuracy of the test set for each benchmark. For METER‑ML, NWPU‑RESISC45, and WHU‑RS19, we adopt the test set splits defined by VHM [vhm]. For UCMerced, we follow the split defined by GeoChat [geochat]. For AID, we present results using both the VHM and GeoChat splits to facilitate fair comparison.
@@ IV-E2 Results
Table IV presents the comparative results across the five scene classification benchmarks. The compared methods include classical VLM baselines (MiniGPTv2 [minigptv2] and LLaVA-1.5 [llava15]), leading open-source VLMs (the QwenVL [qwen25vl], InternVL [internvl25], and MiMo-VL [xiaomi_mimovl] series), and latest RS VLMs (GeoChat [geochat], TEOChat [teochat], LHRS-Bot-Nova [lhrsbot-nova], SkysenseGPT [skysensegpt], VHM [vhm], and ScoreRS [scorers]). As shown, our model consistently surpasses all compared approaches across all benchmarks.
@@ IV-F1 Benchmark and Metric
We evaluate our model’s visual question answering capability using two established benchmarks in the RS domain, including the RSVQA benchmark [rsvqa] and the VQA portion of VRSBench [vrsbench]. The RSVQA compromise two subsets of image-question-answer triplet derived from high‑resolution (HR) orthorectified imagery and low‑resolution (LR) RS data, enabling evaluation of model reasoning across spatial scales. The VRSBench dataset is a large‑scale vision‑language benchmark for RS image understanding that comprises 37,408 question‑answer pairs in test set, supporting a broad range of understanding instructions. The standard question answering accuracy is used as the metric.
@@ IV-F2 Results
Table V presents the results of visual question answering, demonstrating the strong understanding and conversational capabilities of our model. Our approach surpasses all open-source VLMs (including the LLaVA, Qwen, InternVL, and MiMo-VL series) as well as RS VLMs (GeoChat [geochat], LHRS-Bot-Nova [lhrsbot-nova], and VHM [vhm]) across the two benchmarks. In the zero-shot question answering evaluation on VRSBench-VQA [vrsbench], our model outperforms the latest general-purpose models, showing superior generalization ability on RS image question answering.

### v2

@@ IV-D1 Benchmark, Metric, and Comparison Setting
We selected the most widely-used aerial image object detection benchmark, DOTA-v1.0 [63], for our evaluation. The whole DOTA-v1.0 dataset comprises 2,806 high-resolution aerial images and 188,282 object instances across 15 common categories. The proportions of testing set is 1/3. These images were collected from multiple sensors and platforms, and each instance is annotated with a 8 degrees-of-freedom oriented bounding box, capturing the wide variations in object scale, shape, and orientation typical of aerial imagery.
We adopt the Average Precision with no confidence ($\text{AP}_{\text{nc}}$) and report three specific variants: $\text{AP}_{\text{nc}50}$ (IoU threshold is 0.50), $\text{AP}_{\text{nc}75}$ (IoU threshold is 0.75), and $\text{AP}_{\text{nc}50:95}$ (the average $\text{AP}_{\text{nc}}$ computed over IoU thresholds from 0.50 to 0.95 at increments of 0.05). The evaluation is based on the standard MMRotate [79] evaluation procedure. And the splitting length is set 512 with an overlap of 100.
The conventional object detection baselines are trained using the latest MMRotate [79], and the details necessary for reproducibility are also provided in the released code. We obtain a reasonable $\text{AP}_{\text{nc}}$ for comparison methods using the following procedure: We first select a threshold for confidence scores to filter out low-score predictions, and then randomize (or, set to 1) the remaining prediction scores. The AP computed under this condition is denoted as $\text{AP}_{\text{nc}}$ of the conventional detector. To determine an appropriate threshold for each detector, we evaluate $\text{AP}_{\text{nc}}$ on the validation set by varying the confidence threshold from 0.00 to 0.95 in increments of 0.05, and select the threshold that yields the highest $\text{AP}_{\text{nc}}$ for subsequent evaluation on the test set.
@@ IV-D2 Results
To the best of our knowledge, only two existing RS VLMs, LMMRotate [27] (our conference version) and Falcon [21], are capable of performing aerial object detection effectively. However, their common foundation model, Florence-2 [65], employs a fixed and relatively large input size of $\times1024$, which already exceeds the input limit of RSCoVLM. Moreover, LMMRotate is trained specifically for detection, while Falcon performs well only on its training set and does not report test results. In addition, Falcon requires multiple inferences per image, making separate predictions for each category, which results in extremely high computational cost. Therefore, we consider RSCoVLM to be the only vision-language model capable of performing multiple tasks while achieving detection performance that is fairly comparable to specialized object detection models, currently.
@@ IV-E1 Benchmark and Metric
We evaluate our model on five standard remote‑sensing scene‑classification benchmarks. The AID [64] dataset comprises approximately 10,000 images of size $\times600$ pixels across 30 classes. The UCMerced [69] dataset consists of 2,100 images of size $\times256$ pixels covering 21 classes. The NWPU‑RESISC45 [10] dataset contains 31,500 images of size $\times256$ pixels across 45 classes, with large variation in resolution and scene complexity. The WHU‑RS19 [12] dataset includes around 1,000 high‑resolution patches of size $\times600$ pixels spanning 19 classes. The METER‑ML [80] benchmark offers a large‑scale multi‑sensor setup with varied image sizes for extended generalisation evaluation. Together these benchmarks allow a robust assessment of our model’s generalisation across dataset scale, class‑set size, imaging conditions and spatial resolutions.
We report overall accuracy of the test set for each benchmark. For METER‑ML, NWPU‑RESISC45, and WHU‑RS19, we adopt the test set splits defined by VHM [46]. For UCMerced, we follow the split defined by GeoChat [22]. For AID, we present results using both the VHM and GeoChat splits to facilitate fair comparison.
@@ IV-E2 Results
Table IV presents the comparative results across the five scene classification benchmarks. The compared methods include classical VLM baselines (MiniGPTv2 [6] and LLaVA-1.5 [37]), leading open-source VLMs (the QwenVL [2], InternVL [8], and MiMo-VL [70] series), and latest RS VLMs (GeoChat [22], TEOChat [20], LHRS-Bot-Nova [34], SkysenseGPT [41], VHM [46], and ScoreRS [44]). As shown, our model consistently surpasses all compared approaches across all benchmarks.
@@ IV-F1 Benchmark and Metric
We evaluate our model’s visual question answering capability using two established benchmarks in the RS domain, including the RSVQA benchmark [39] and the VQA portion of VRSBench [30]. The RSVQA compromise two subsets of image-question-answer triplet derived from high‑resolution (HR) orthorectified imagery and low‑resolution (LR) RS data, enabling evaluation of model reasoning across spatial scales. The VRSBench dataset is a large‑scale vision‑language benchmark for RS image understanding that comprises 37,408 question‑answer pairs in test set, supporting a broad range of understanding instructions. The standard question answering accuracy is used as the metric.
@@ IV-F2 Results
Table V presents the results of visual question answering, demonstrating the strong understanding and conversational capabilities of our model. Our approach surpasses all open-source VLMs (including the LLaVA, Qwen, InternVL, and MiMo-VL series) as well as RS VLMs (GeoChat [22], LHRS-Bot-Nova [34], and VHM [46]) across the two benchmarks. In the zero-shot question answering evaluation on VRSBench-VQA [30], our model outperforms the latest general-purpose models, showing superior generalization ability on RS image question answering.